# Lumen: Unleashing Versatile Vision-Centric Capabilities of Large Multimodal Models

### v1

@@ Abstract
Large Multimodal Model (LMM) is a hot research topic in the computer vision area and has also demonstrated remarkable potential across multiple disciplinary fields. A recent trend is to further extend and enhance the perception capabilities of LMMs. The current methods follow the paradigm of adapting the visual task outputs to the format of the language model, which is the main component of a LMM. This adaptation leads to convenient development of such LMMs with minimal modifications, however, it overlooks the intrinsic characteristics of diverse visual tasks and hinders the learning of perception capabilities. To address this issue, we propose a novel LMM architecture named Lumen, a Large multimodal model with versatile vision-centric capability enhancement. We decouple the LMM’s learning of perception capabilities into task-agnostic and task-specific stages. Lumen first promotes fine-grained vision-language concept alignment, which is the fundamental capability for various visual tasks. Thus the output of the task-agnostic stage is a shared representation for all the tasks we address in this paper. Then the task-specific decoding is carried out by flexibly routing the shared representation to lightweight task decoders with negligible training efforts. Benefiting from such a decoupled design, our Lumen surpasses existing LMM-based approaches on the COCO detection benchmark with a clear margin and exhibits seamless scalability to additional visual tasks. Furthermore, we also conduct comprehensive ablation studies and generalization evaluations for deeper insights. The code will be released at https://github.com/SxJyJay/Lumen.
Keywords:
Large multimodal models, Vision-centric capabilities
@@ 1 Introduction
While the Large Language Models (LLMs) are currently kindling the spark of Artificial General Intelligence (AGI), Large Multimodal Models (LMMs) [69, 35, 50, 53] take a step forward by integrating visual modalities with the linguistic prowess of LLMs. With the instruction-following and content-reasoning capabilities inherited from LLMs, the LMM has successfully functioned as a versatile assistant across a wide range of tasks, including visual question answering [2, 40], image captioning [47, 21], visual commonsense reasoning [3, 37], etc.
@@
Figure 1: The comparison of our proposed Lumen with previous methods (e.g., Griffon [65]). Previous methods serialize bounding box coordinates into discrete token sequences to conform to the language-oriented outputs format of LMMs, disregarding the unordered nature inherent in bounding boxes. Our Lumen first predicts unified heatmaps for various tasks. These heatmaps are further used for guiding simple decoding tools with the parsed task-type indicators to support versatile visual tasks. We omit the user instruction of referring to all persons in the image for conciseness.
@@
Although convenient, the aforementioned approaches encounter challenges when scaling up to more intricate scenarios and vision-centric tasks. First, those LLMs highly relies on language models’s auto-regressive sequence generation method, which leads to high uncertainty when multiple objects are concurrently addressed. As shown in the first row of Fig 1, the three persons highlighted with bounding boxes of different colors lack an inherent order. And imposing a sequential order on them would exacerbate the confusion in the decoding process, as the model would be compelled to produce drastically different outputs after the same word “person”. Although sequence augmentation techniques introduced in Pix2Seq [10] can alleviate this issue, they are tailored for the object detection task and can not be seamlessly transferred to other vision-centric tasks such as instance segmentation and pose estimation. On the other hand, in contrast to tasks within the NLP field, which often exhibit stronger inter-task correlations [56], vision-centric tasks are inherently discrete due to variations in the granularity of visual content comprehension [59]. Therefore, reformulating heterogeneous vision-centric task outputs into language-oriented formats tends to overemphasize the format consistency with language models, while the learning of the underlying visual perception capabilities and intrinsic characteristics of diverse vision tasks is overlooked.
@@
When delving into fundamental vision-centric tasks, namely object detection, instance segmentation, and pose estimation [^1] We only investigate these three visual tasks in this paper, as LMMs inherently lack image generation capabilities., we observe that they can be decoupled into task-agnostic and task-specific learning processes. Essentially, these tasks share a common task-agnostic objective of identifying individual instances with an instruction like “finding the area addressed in instructions”, while their task-specific definitions introduce different decoding rules for generating diverse formats of outputs (e.g., boxes, masks, or points). Compared with the general task-agnostic objective, task-specific outputs relies less on semantics but is more difficult for the LMMs to learn to output. Based on the above analysis, in this paper, we propose Lumen, a Large multimodal model with vision-centric capabilities enhancement, which decouples the task-agnostic and task-specific learning in two consecutive stages as shown in the Fig 1. Concretely, in the first stage, the aforementioned visual tasks are unified by reformulating them into the same matching problem. We start by feeding the user’s instruction and image into a LMM for content comprehension. The obtained responses contain a designed special token (i.e, [LOC] in Fig 1) that encapsulates the visual concepts conveyed in the provided instruction, regardless of the specific task. Subsequently, this special token interacts with image patches via a transformer-based aligner to generate a heatmap, wherein the response at each location indicates the matching probability between the instruction and the corresponding image region. In the second stage, utilizing this heatmap as indicative guidance, task-specific decoding processes are further managed by flexibly assembling predefined decoding rules and lightweight decoders to generate the final outputs with different formats. With such decoupled learning behaviors, our Lumen can concentrate on promoting fine-grained multimodal content comprehension, rather than being trapped in learning diverse specialized decoding rules lacking in semantics.
@@
In summary, our contributions are three folds: (1) We introduce Lumen, a Large multimodal model with vision-centric capabilities enhancement, which decouples the task-agnostic and task-specific learning processes for efficiently unleashing the vision-centric potential of the LMM; (2) Our Lumen can seamlessly adapt to tasks such as object detection, instance segmentation, and pose estimation without requiring specialized dialogue datasets as done in the previous work [65]. (3) Our Lumen not only significantly outperforms existing LMM-based competitors in challenging object detection by a significant margin, but also achieves comparable performances to specialist models on several tasks, demonstrating the effectiveness of our method. Additionally, we conduct extensive ablation studies and generalization evaluations for further in-depth analysis.
@@ 2.2 Vision Generalist Models
Generalist models [52, 51, 49, 36] in the vision domain aim at unifying a wide range of vision tasks using one model. Motivated by the success of sequence-to-sequence models in NLP field [47], OFA [54] and GIT [49] unify various tasks in the sequence generation format. Following this trend, Unified-IO [36], Pix2Seq v2 [11] and UniTab [63] add discrete coordinate tokens into the vocabulary, so as to accommodating more tasks. Moreover, Gato [43] successfully unifies reinforcement learning tasks as the sequence generation format. Nevertheless, the sequence generation modeling can lead to low inference speeds and degraded performances. Toward this end, Uni-Perceivers [70, 29] unify different tasks as the maximum likelihood matching problem by calculating representation similarities of all targets and each input. With such an objective, generative and non-generative tasks can be unified by selecting corresponding input and target candidates. However, these generalist models are restricted to pre-defined tasks, failing to be flexibly instructed by natural languages like LMMs.
@@ 2.3 LMMs with Vision-Centric Capabilities
Figure 2: The overall framework of the proposed Lumen. Our Lumen consists of two stages. In the first stage, the input image and the instruction with designed special tokens are embedded and fed into a large language model to interact and comprehend visual and textual contents. Then, the [LOC] token output and high-resolution image features are further aligned to produce a heatmap denoting cross-modal matching probabilities. In the second stage, the heatmap can serve as a strong indication for various vision tasks, and the task outputs can be obtained with lightweight task-specific decoders. The routing of the decoding pathway is determined by the task token (e.g., [DET] in image) output generated in the first stage.
@@ 3.1 Stage 1: Task-Agnostic Matching
3.1.1 Conversation Reformulation.
A preliminary step for adapting vision-centric tasks to LMMs is to reformulate the visual data into conversation formats. For different tasks, we employ a unified instruction-response template: “USER: [IMG]. Please find the location of {description}. Respond with {format}. ASSISTANT: Sure, the location is [LOC]. The task output is [DET]/[SEG]/...” Here, {description} and {format} can be customized according to specific tasks. For vision-centric tasks like object detection, instance segmentation and pose estimation, {description} is a certain class name, and {format} can be boxes, masks, or points. For vision-language tasks like visual grounding and referring segmentation, {description} is the referring sentence and {format} can be boxes or masks. Apart from [IMG] that denotes image features, we also introduce a task-agnostic special token [LOC] and a set of task-specific special tokens, including [DET], [SEG], [POINT], [GROUND] and [REFSEG]. As shown in Fig 2, regardless of task types, [LOC] is tasked to summarize the visual concepts in instructions and further densely align with image regions. Task-specific tokens serve as routers for guiding the decoding pathway in the second stage. Examples of reformulated conversation data of various tasks are illustrated in Fig 3.
Figure 3: The illustration of conversation data examples reformulated from different vision-centric tasks and vision-language tasks, as well as the data sample of the traditional VQA task for comparison. For each reformulated data example, we sequentially present the original image (left), the heatmap generated from annotations (middle), and the task-specific ground truth (right) in the figure.
3.1.2 Model Architecture.
With the given image $x_{img}$ and reformulated instruction $x_{txt}$ as inputs, we first feed them into a conventional LMM (LLaVA-1.1 [35] in our implementation) to generate language responses. We also extract high resolution image features $f_{img}\in\mathbb{R}^{32\times 32\times 256}$ following prior works [4, 65]:
@@
where $\mathcal{F_{V}}$ denotes the vision encoder. Since the adopted vision encoder, CLIP-ViT [41], is pretrained to accept image inputs of $224\times 224$ pixels, we resize its positional embeddings to adapt to higher resolution, e.g., $448\times 448$ pixels. Afterward, the feature of [LOC] token $h_{loc}\in\mathbb{R}^{1\times 256}$ from the generated response, together with high resolution image features, are fed into the proposed V-L dense aligner to calculate their matching probabilities, resulting in a heatmap $\hat{H}\in\mathbb{R}^{32\times 32}$:
@@
where $\mathcal{F_{A}}$ represents the V-L dense aligner. In our implementation, we employ a lightweight transformer-based architecture as $\mathcal{F_{A}}$ to interact high resolution image feature tokens and the [LOC] token. More discussions on the selection of architecture design are included in Sec.4.3.
3.1.3 Model Training.
To enrich semantics during the dense alignment, we merge data from various tasks by reformulating them into uniform conversation formats as shown in Fig 3. For various tasks, we hide diverse task-specific output format details and their learning targets are reformulated and unified as heatmaps, where each element denotes the matching probability between the instruction and the corresponding region. To construct the ground-truth heatmap $H$, we use a Gaussian kernel to fill the probabilities into a blank map following prior works [67, 48]:
@@ 3.2 Stage 2: Task-Specific Decoding
3.2.1 Decoding Modules.
We first introduce the operations of the three key modules illustrated in Fig 2. (1) Peak point selection parses the heatmap into a set of points, where each point indicates the center location of an identified object or keypoint. Specifically, we filter the locations with heatmap response greater than their 8 surrounding neighbors as candidate peak points, and then retain the top $K$ candidates as the selected results. The value of $K$ can vary across different tasks, which will be elaborated on in Sec.4.1. (2) Box decoder is used for further regressing the extent of the objects designated by the selected peak points. For efficiency, instead of introducing an additional network as the box decoder, we reuse the V-L dense aligner by appending two new learnable special tokens after the [LOC] token as additional text-side inputs. Accordingly, the V-L dense aligner will generate two additional 1-channel map predictions, which are used for regressing the height and width under the supervision of the L1 loss functions similar to [67]. (3) Promptable mask decoder can accept both points and boxes as visual instructions to generate the mask for the referred object. We directly use the SAM model [25] for executing the above process without finetuning. In general, these decoding modules introduce negligible training costs and can be easily implemented.
3.2.2 Decoding Pathways.
On the basis of these three decoding modules, we customize their different cooperation patterns to perform diverse tasks, which will be introduced as follows: (1) Pose estimation requires to return the keypoint coordinates. We can easily obtain these coordinates by parsing the predicted heatmap with the peak point selection operation. (2) Object detection and visual grounding share the same output format of bounding box, therefore, we employ the identical decoding pathway for them. Specifically, as shown in Fig 2, we feed the heatmap into the cascaded peak point selection and box decoder modules to generate bounding boxes. (3) Instance segmentation and referring segmentation share a relationship analogous to the one between object detection and visual grounding, therefore, we also adopt the same decoding pathway for them. Concretely, we first parse both the center point and bounding box of the object, and then, we use them as visual instructions to guide the promptable mask decoder to generate the mask prediction.
@@ 4.1 Experimental Setup
Our training data consists of datasets from the following different tasks. (1) For object detection, we use MSCOCO [32], Objects365 [45] and OpenImages [27]. (2) For visual grounding, we use RefCOCO, RefCCO+ and RefCOCOg [64]. (3) For pose estimation, we use MSCOCO keypoints [32] and AIC [57]. (4) For visual question answering, we use LLaVA-Instruct-150k dataset [35]. It is worth noting that for instance segmentation and referring segmentation, since we first transform their masks into bounding boxes as mentioned in Sec 3.1.2, the constructed heatmaps are the same as the object detection. Additionally, given that we do not finetune the mask decoder in the second decoding stage, we actually do not use segmentation annotations throughout the entire training process.
4.1.2 Model Configurations.
For the first task-agnostic matching stage, we utilize pre-trained CLIP ViT-L/14 [41] and Vicuna-7B [12] as our vision encoder and large language model, respectively.
@@
For the second decoding stage, since the different decoding modules and pathways have been introduced in Sec 3.2, we primarily specify the $K$ value choices and corresponding post-processing details of different tasks here. (1) For object detection and instance segmentation, which generally includes multiple objects in a given image, we set $K=100$ to generate 100 box candidates. Then we apply regular NMS [18] to filter redundant boxes, and the remaining boxes are used for further prompt mask decoder to generate instance segmentation results. (2) For visual grounding and referring segmentation, given that only one object is referred by a referring sentence, we set $K=1$ to only generate one predicted box/mask, and no post-processing is required. (3) For pose estimation, we follow previous works [16, 10] to first crop the single object from the image using bounding boxes. Then, for each keypoint category, we set $K=1$ to extract the point with the highest matching probability as the prediction result.
4.1.3 Training Details.
@@
For object detection, visual grounding, pose estimation and visual question answering, we set their sampling rates as 0.56, 0.19, 0.19 and 0.06, respectively, for balanced data sampling. We set the batch size to 20 and train the model for 50,000 steps on 8 NVIDIA 80G A100 GPUs. The loss function balance terms $\lambda_{h}$ and $\lambda_{t}$ are both set to 1. We use AdamW as the optimizer with an initial learning rate of $3\times 10^{4}$ and weight decay of 0. During training, we do not calculate heatmap loss $\mathcal{L}_{h}$ for visual question answering data. We do not use any data augmentation techniques for all tasks for simplicity.
4.1.4 Evaluation Metrics.
[…]
Table 1: Results on fundamental vision-centric tasks and vision-language tasks. Methods in “Vision Generalists” adopt strong visual backbones and additional designs adaptive to vision-centric tasks. On the contrary, methods in “LMM Generalists" only inherit the LMM’s visual backbones, which are proficient in multimodal comprehension but weak in capturing detailed object features. This line of methods aligns with the pursuit of unleashing the inherent vision-centric capabilities of LMMs, rather than relying on external visual experts as proxies. We use $\dagger$ to indicate that the visual capabilities of VisionLLM stem from external vision encoders. “-” indicates the result is not reported or the corresponding task is not supported by the method.
@@
Faster R-CNN [44] 40.3 61.0 44.0 - - - - - - - -
DETR [5] 43.3 63.1 45.9 - - - - - - - -
Mask R-CNN [17] 41.0 61.7 44.9 37.1 58.4 40.1 - - - - -
PolarMask [58] - - - 30.5 52.0 31.1 - - - - -
CPM [58] - - - - - - 62.7 86.2 70.9 - -
RTMPose [20] - - - - - - 68.2 88.3 75.9 - -
MDETR [24] - - - - - - - - - 83.4 -
SeqTR [68] - - - - - - - - - 82.7 64.7
LAVT [62] - - - - - - - - - - 61.2
ReLA [33] - - - - - - - - - - 65.0
@@
Pix2Seq-v2 [10] 46.5 - - 38.2 - - 64.8 - - - -
Uni-Perceiver-v2 [29] 58.6 - - 50.6 - - - - - - -
mPLUG-2 [60] 46.9 - - 40.6 - - - - - 84.7 -
${}^{\dagger}$ [55] 44.6 64.0 48.1 25.1 50.0 22.4 - - - - -
@@
Shikra-7B [8] - - - - - - - - - 82.3 -
MiniGPT-v2-7B [6] - - - - - - - - - 84.4 -
Griffon-13B [65] 23.2 37.6 23.4 - - - - - - 81.5 -
InstructCV [15] - 48.5 - - - - - - - - -
Lumen-7B (Ours) 33.9 51.2 34.2 29.1 47.5 29.6 65.4 90.4 72.2 83.5 64.0

### v2

@@ Abstract
Large Multimodal Model (LMM) is a hot research topic in the computer vision area and has also demonstrated remarkable potential across multiple disciplinary fields. A recent trend is to further extend and enhance the perception capabilities of LMMs. The current methods follow the paradigm of adapting the visual task outputs to language-oriented formats. This adaptation leads to the convenient development of such LMMs with minimal modifications, however, it overlooks the intrinsic characteristics of diverse visual tasks and hinders the learning of perception capabilities. To address this issue, we propose a novel LMM architecture named Lumen, which decouples the learning of perception capabilities into task-agnostic and task-specific stages. Firstly, Lumen promotes fine-grained vision-language concept alignment, which is the fundamental capability for various visual tasks. Thus the output of the task-agnostic stage is a shared representation for all vision-centric tasks we address in this paper. Afterward, the task-specific decoding is carried out by flexibly routing the shared representation to lightweight task decoders with negligible training efforts. Comprehensive experimental results on a series of vision-centric and VQA benchmarks indicate that our Lumen model not only achieves or surpasses the performance of existing LMM-based approaches in a range of vision-centric tasks while maintaining general visual understanding and instruction following capabilities.
@@ 1 Introduction
As the Large Language Models (LLMs) currently kindle the spark of Artificial General Intelligence (AGI), Large Multimodal Models (LMMs) [1, 2] take a step forward by integrating visual modalities with the linguistic prowess of LLMs. With the instruction-following and content-reasoning capabilities inherited from LLMs, the LMM has successfully functioned as a versatile assistant across a wide range of tasks, including visual question answering [3, 4], image captioning [5, 6], visual commonsense reasoning [7, 8], etc.
@@
Figure 1: Comparison of our proposed Lumen with previous methods. Our Lumen first predicts unified heatmaps for various tasks. These heatmaps are further used for guiding simple decoding tools with the parsed task-type indicators to support versatile visual tasks. We omit the user instruction of referring to all persons in the image for conciseness.
@@
Although convenient, the aforementioned approaches encounter challenges when scaling up to more intricate scenarios and vision-centric tasks. First, those LLMs highly relies on language models’s auto-regressive sequence generation method, which leads to high uncertainty when multiple objects are concurrently addressed. As shown in the first row of Fig 1, the three persons highlighted with bounding boxes of different colors lack an inherent order. And imposing a sequential order on them would exacerbate the confusion in the decoding process, as the model would be compelled to produce drastically different outputs after the same word “person”. Although sequence augmentation techniques introduced in Pix2Seq [13] can alleviate this issue, they are tailored for the object detection task and can not be seamlessly transferred to other vision-centric tasks such as instance segmentation and pose estimation. On the other hand, in contrast to tasks within the NLP field, which often exhibit stronger inter-task correlations [14], vision-centric tasks are inherently discrete due to the inductive bias introduced by their distinct task definitions [15]. Therefore, reformulating heterogeneous vision-centric task outputs into language-oriented formats tends to overemphasize the format consistency with language models, while the learning of the underlying visual perception capabilities and intrinsic characteristics of diverse vision tasks is overlooked.
@@
When delving into fundamental vision-centric tasks, namely object detection, instance segmentation, and pose estimation [^1] We only investigate these three visual tasks in this paper, as LMMs inherently lack image generation capabilities., we observe that they can be decoupled into task-agnostic and task-specific learning processes. Essentially, these tasks share a common task-agnostic objective of identifying individual instances with an instruction like “finding the area addressed in instructions”, while their task-specific definitions introduce different decoding rules for generating diverse formats of outputs (e.g., boxes, masks, or points). Compared with the general task-agnostic objective, task-specific outputs relies less on semantics but is more difficult for the LMMs to learn to output. Based on the above analysis, in this paper, we propose Lumen, a Large multimodal model with vision-centric capabilities enhancement, which decouples the task-agnostic and task-specific learning in two consecutive stages as shown in the Fig 1. Concretely, in the first stage, the aforementioned visual tasks are unified by reformulating them into the same matching problem. We start by feeding the user’s instruction and image into a LMM for content comprehension. The obtained responses contain a designed special token (i.e, [LOC] in Fig 1) that encapsulates the visual concepts conveyed in the provided instruction, regardless of the specific task. Subsequently, this special token interacts with image patches via a transformer-based aligner to generate a heatmap, wherein the response at each location indicates the matching probability between the instruction and the corresponding image region. In the second stage, utilizing this heatmap as indicative guidance, task-specific decoding processes are further managed by flexibly assembling predefined decoding rules and lightweight decoders to generate the final outputs with different formats. Due to such decoupled learning behaviors, on the one hand, Lumen can concentrate on promoting fine-grained multimodal content comprehension, rather than being trapped in learning diverse specialized decoding rules lacking in semantics. On the other hand, Lumen can be affected less by the various inductive biases associated with vision-centric tasks within the LLM token space, thereby seamlessly maintaining general visual understanding and instruction following capabilities, as demonstrated in Table 2.
@@
In summary, our contributions are three folds: (1) We introduce Lumen, a Large multimodal model with vision-centric capabilities enhancement, which unleashes the vision-centric potential of the LMM by decoupling the task-agnostic and task-specific learning processes; (2) Our Lumen can seamlessly adapt to tasks such as object detection, instance segmentation, and pose estimation without requiring specialized dialogue datasets as done in the previous work [12]. (3) Our Lumen not only matches or exceeds the performances of existing LMM-based approaches on a series of vision-centric tasks, but also maintains general visual understanding and instruction following capabilities.
@@ 2.2 Vision Generalist Models
Generalist models in the vision domain aim at unifying a wide range of vision tasks using one model. Motivated by the success of sequence-to-sequence models in NLP field [5], OFA [20] and GIT [21] unify various tasks in the sequence generation format. Following this trend, Unified-IO [22], Pix2Seq v2 [23] and UniTab [24] add discrete coordinate tokens into the vocabulary, so as to accommodating more tasks. Moreover, Gato [25] successfully unifies reinforcement learning tasks as the sequence generation format. Nevertheless, the sequence generation modeling can lead to low inference speeds and degraded performances. Toward this end, Uni-Perceivers [26, 27] unify different tasks as the maximum likelihood matching problem by calculating representation similarities of all targets and each input. With such an objective, generative and non-generative tasks can be unified by selecting corresponding input and target candidates. However, these generalist models are restricted to pre-defined tasks, failing to be flexibly instructed by natural languages like LMMs.
@@ 2.3 LMMs with Vision-Centric Capabilities
Figure 2: Overall framework of the proposed Lumen. Our Lumen consists of two stages. In the first stage, the input image and the instruction with designed special tokens are embedded and fed into a large language model to interact and comprehend visual and textual contents. Then, the [LOC] token output and high-resolution image features are further aligned to produce a heatmap denoting cross-modal matching probabilities. In the second stage, the heatmap can serve as a strong indication for various vision tasks, and the task outputs can be obtained with lightweight task-specific decoders. The routing of the decoding pathway is determined by the task token (e.g., [DET] in image) output generated in the first stage.
@@ 3.1 Stage 1: Task-Agnostic Matching
3.1.1 Conversation Reformulation
A preliminary step for adapting vision-centric tasks to LMMs is to reformulate the visual data into conversation formats. For different tasks, we employ a unified instruction-response template: “USER: [IMG]. Please find the location of {description}. Respond with {format}. ASSISTANT: Sure, the location is [LOC]. The task output is [DET]/[SEG]/...” Here, {description} and {format} can be customized according to specific tasks. For vision-centric tasks like object detection, instance segmentation and pose estimation, {description} is a certain class name, and {format} can be boxes, masks, or points. For vision-language tasks like visual grounding and referring segmentation, {description} is the referring sentence and {format} can be boxes or masks. Apart from [IMG] that denotes image features, we also introduce a task-agnostic special token [LOC] and a set of task-specific special tokens, including [DET], [SEG], [POINT], [GROUND] and [REFSEG]. As shown in Fig 2, regardless of task types, [LOC] is tasked to summarize the visual concepts in instructions and further densely align with image regions. Task-specific tokens merely serve as routers for guiding the decoding pathway in the second stage without encoding task-specific inductive bias. Examples of reformulated conversation data of various tasks are illustrated in Fig 3.
Figure 3: The illustration of reformulated conversation data examples of different tasks. For each reformulated data example, we sequentially present the original image (left), the heatmap generated from annotations (middle), and the task-specific ground truth (right) in the figure.
3.1.2 Model Architecture
With the given image $x_{img}$ and reformulated instruction $x_{txt}$ as inputs, we first feed them into a conventional LMM (LLaVA-1.5 [38] in our implementation) to generate language responses. We also extract high resolution image features $f_{img}\in\mathbb{R}^{32\times 32\times 256}$ following prior works [11, 12]:
@@ 3.1.2 Model Architecture
where $\mathcal{F_{V}}$ denotes the vision encoder. Since the adopted vision encoder, CLIP-ViT [39], is pretrained to accept image inputs of $336\times 336$ pixels, we resize its positional embeddings to adapt to higher resolution, e.g., $448\times 448$ pixels. Afterward, the feature of [LOC] token $h_{loc}\in\mathbb{R}^{1\times 256}$ from the generated response, together with high-resolution image features, are fed into the proposed V-L dense aligner to calculate their matching probabilities, resulting in a heatmap $\hat{H}\in\mathbb{R}^{32\times 32}$:
@@
where $\mathcal{F_{A}}$ represents the V-L dense aligner. In our implementation, we employ a lightweight transformer-based architecture as $\mathcal{F_{A}}$ to interact high resolution image feature tokens and the [LOC] token. More discussions on the architecture design are included in Sec.4.3 and Appendix A.1.
3.1.3 Model Training
To enrich semantics during the dense alignment, we merge data from various tasks by reformulating them into uniform conversation formats as shown in Fig 3. For vision-centric tasks, we hide diverse task-specific output format details and their learning targets are reformulated and unified as heatmaps, where each element denotes the matching probability between the instruction and the corresponding region. To construct the ground-truth heatmap $H$, we use a Gaussian kernel to fill the probabilities into a blank map following prior works [40, 41]:
@@ 3.2 Stage 2: Task-Specific Decoding
Decoding Modules. We first introduce the operations of the three key modules illustrated in Fig 2. (1) Peak point selection parses the heatmap into a set of points, where each point indicates the center location of an identified object or keypoint. Specifically, we filter the locations with heatmap response greater than their 8 surrounding neighbors as candidate peak points, and then retain the top $K$ candidates as the selected results. The value of $K$ can vary across different tasks, which will be elaborated on in Sec.4.1. (2) Box decoder is used for further regressing the extent of the objects designated by the selected peak points. For efficiency, instead of introducing an additional network as the box decoder, we reuse the V-L dense aligner by appending two new learnable special tokens after the [LOC] token as additional text-side inputs. Accordingly, the V-L dense aligner will generate two additional 1-channel map predictions, which are used for regressing the height and width under the supervision of the L1 loss functions similar to [40]. (3) Promptable mask decoder can accept both points and boxes as visual instructions to generate the mask for the referred object. We directly use the SAM model [34] for executing the above process without finetuning. In general, these decoding modules introduce negligible training costs and can be easily implemented.
Decoding Pathways. On the basis of these three decoding modules, we customize their different cooperation patterns to perform diverse tasks, which will be introduced as follows: (1) Pose estimation requires to return the keypoint coordinates. We can easily obtain these coordinates by parsing the predicted heatmap with the peak point selection operation. (2) Object detection and visual grounding share the same output format of bounding box, therefore, we employ the identical decoding pathway for them. Specifically, as shown in Fig 2, we feed the heatmap into the cascaded peak point selection and box decoder modules to generate bounding boxes. (3) Instance segmentation and referring segmentation share a relationship analogous to the one between object detection and visual grounding, therefore, we also adopt the same decoding pathway for them. Concretely, we first parse both the center point and bounding box of the object, and then, we use them as visual instructions to guide the promptable mask decoder to generate the mask prediction.
@@ 4.1 Experimental Setup
Datasets. Our training data consists of datasets from the following different tasks. (1) For object detection, we use MSCOCO [43], Objects365 [44] and OpenImages [45]. (2) For visual grounding, we use RefCOCO, RefCCO+ and RefCOCOg [46]. (3) For pose estimation, we use MSCOCO keypoints [43] and AIC [47]. (4) For visual question answering, we employ a subset of ShareGPT4V dataset [48] with 665K samples. It is worth noting that for instance segmentation and referring segmentation, since we first transform their masks into bounding boxes as mentioned in Sec 3.1.2, the constructed heatmaps are the same as the object detection. Additionally, given that we do not finetune the mask decoder in the second decoding stage, we actually do not use segmentation annotations throughout the entire training process.
@@
Model Configurations.  For the first task-agnostic matching stage, we utilize pre-trained CLIP ViT-L/14-336 [39] and LLaVA-7B-v1.5 [2] as our vision encoder and large multimodal model, respectively. For the second decoding stage, since the different decoding modules and pathways have been introduced in Sec 3.2, we primarily specify the $K$ value choices and corresponding post-processing details of different tasks here. (1) For object detection and instance segmentation, which generally includes multiple objects in a given image, we set $K=100$ to generate 100 box candidates. Then we apply regular NMS [49] to filter redundant boxes, and the remaining boxes are used for further prompt mask decoder to generate instance segmentation results. (2) For visual grounding and referring segmentation, given that only one object is referred by a referring sentence, we set $K=1$ to only generate one predicted box/mask, and no post-processing is required. (3) For pose estimation, we follow previous works [50, 13] to first crop the single object from the image using bounding boxes. Then, for each keypoint category, we set $K=1$ to extract the point with the highest matching probability as the prediction result.
@@
Training Details.  For the task-agnostic stage, our training comprises two phases. (1) In Phase 1, we mix the object detection, visual grounding and pose estimation data with sampling rates of 0.69, 0.23, 0.08, respectively, for balanced data sampling. (2) In Phase 2, we mix the visual question-answering, object detection, visual grounding and pose estimation data with sample rates of 0.67, 0.23, 0.07, 0.03, respectively. We set the batch size to 160 and train the first step for 50,000 steps and the second step for 10,000 steps on 8 NVIDIA 80G A100 GPUs. The loss function balance terms $\lambda_{h}$ and $\lambda_{t}$ are both set to 1. For each phase, we use AdamW as the optimizer with an initial learning rate of $3\times 10^{-4}$ and weight decay of 0. During training, we do not calculate heatmap loss $\mathcal{L}_{h}$ for visual question-answering data. We do not use any data augmentation techniques for all tasks.
Evaluation Metrics. We adopt evaluation metrics commonly used within each field of task. For object detection and instance segmentation, we use mAP based on box IoU and mask IoU, respectively. For pose estimation, we use mAP based on OKS (object keypoint similarity). For visual question-answering, we comply with the evaluation protocol of each individual benchmark.
Table 1: Results on fundamental vision-centric tasks and vision-language tasks. We use “-” and gray cell to indicate the result is not reported and the corresponding task is not supported by the method, respectively.
@@
Faster R-CNN [51] 40.3 61.0 44.0
DETR [52] 43.3 63.1 45.9
Mask R-CNN [53] 41.0 61.7 44.9 37.1 58.4 40.1
PolarMask [54] 30.5 52.0 31.1
CPM [54] 62.7 86.2 70.9
RTMPose [55] 68.2 88.3 75.9
MDETR [56] 83.4
SeqTR [57] 82.7 64.7
LAVT [58] 61.2
ReLA [59] 65.0
@@
Pix2Seq-v2 [13] 46.5 - - 38.2 - - 64.8 - -
Uni-Perceiver-v2 [27] 58.6 - - 50.6 - -
mPLUG-2 [60] 46.9 - - 40.6 - - 84.7
VisionLLM [36] 44.6 64.0 48.1 25.1 50.0 22.4 -
@@
Shikra-7B [31] 82.3
MiniGPT-v2-7B [9] 84.4
Griffon-13B [12] 23.2 37.6 23.4 81.5
InstructCV [61] - 48.5 -
Lumen-7B (Ours) 35.3 53.2 35.8 30.4 49.8 31.0 67.2 90.4 75.6 83.6 65.1
Table 2: Results on prevalent VQA benchmarks. Here, we employ English MMBench dev, SEEDBench image, MME test, MMMU val and MathVista mini sets for evaluation.
Method Param MMBench SEED MME MMMU MathVista
InstructBLIP [19] 7B 36.0 58.8 1213/292 32.9 25.3
MiniGPT-4 [1] 7B 24.3 47.4 582/144 - 23.1
Shikra [31] 7B 58.8 - - - -
Qwen-VL-Chat [11] 7B 60.6 58.2 1488/361 35.9 -
LLaVA-v1.5 [38] 7B 64.3 66.1 1511/296 35.6 23.5
Lumen (Ours) 7B 64.9 65.8 1426/332 35.2 24.6